1. Calculate the break-even point for a new line of shirts. Selling price will be $48 per shirt. Labor costs will be $12 per shirt. Administrative costs allocated to the shirt line are estimated to be $80,000 annually and the sales and marketing expenses are $40,000 a year for the shirt line. Cost of materials will be $20 per shirt. What is the break-even quantity? The break-even quantity lies in the range:

-    more than 8,000 shirts.

-    *6,001 to 8,000 shirts.

-    3,500 to 6,000 shirts.

-    fewer than 3,500 shirts.
Answer #1

Break even quantity = Total fixed costs/contribution margin per unit

Total fixed costs = 80,000+40,000 = $120,000

Contribution Margin per unit = Selling price per unit - variable cost per unit

= 48-12-20

= $16 per unit

Hence, Break even quantity = 120,000/16

= 7,500 shirts

hence, the answer is

6,001 to 8,000 shirts.